A solution contains 8.37×10-3 M potassium fluoride and 7.91×10-3 M sodium carbonate.  
Solid magnesium acetate is added slowly to this mixture.

A. What is the formula of the substance that precipitates first?

formula =




B. What is the concentration of magnesium ion when this precipitation first begins?  
[Mg2+] =  M
Answer #1

A) Ksp of MgCO3 = 4.0 x 10-5

MgCO3 .............> Mg^2+ + CO3^2-

Ksp = [Mg^2+] [CO3^2-]

4.0 x 10-5 = s * s

or

s = 6.32 * 10^-3 mole / L

thus

molar solubility of MgCO3 = 6.32 * 10^-3 mole / L

and

Ksp of MgF2 = 6.4 x 10-9

MgF2 ..........> Mg^2+ + 2 F-

.........................s............2s

Ksp = s * (2s)^2 = 4s^3 = 6.4 x 10-9

s = 4.0 * 10^-5 M

molar solubility of MgF2 = 4.0 * 10^-5 M

thus

First MgF2 precipitated.

Formula is MgF2.

B)

MgF2 ..........> Mg^2+ + 2 F-

Ksp = [Mg^2+][F-]^2

6.4 x 10-9 = [Mg^2+] * (8.37×10-3)^2

or

[Mg^2+] = 9.14 * 10^-5 M

concentration of magnesium ion = 9.14 * 10^-5 M